### v1

@@ 1 Preliminaries: the algebraic classification
Theorem 8 (see [2]).
Let $\mathfrak{B}$ be a 4-dimensional complex nilpotent Bol algebra. Then $\mathfrak{B}$ is isomorphic to one of these Lie-Yamaguti algebras $\mathcal{L}_{01},\dots,\mathcal{L}_{41},$ $\mathcal{L}_{45}^{\alpha},\dots,\mathcal{L}_{52}$ or one of the following non-mutually isomorphic algebras:
$\mathfrak{B}_{01}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=-\frac{1}{2}e_{4},$
$[e_{1},e_{3},e_{2}]=\frac{1}{2}e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathfrak{B}_{02}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=-\frac{1}{2}e_{4},$
$[e_{1},e_{3},e_{2}]=\frac{1}{2}e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathfrak{B}_{03}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=-\frac{1}{2}e_{4},$
$[e_{1},e_{3},e_{2}]=\frac{1}{2}e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathfrak{B}_{04}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=(\alpha-1)e_{4},$ $[e_{1},e_{3},e_{2}]=\alpha e_{4}.$
$\mathfrak{B}_{05}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{3}]=e_{4},$
$[e_{2},e_{3},e_{1}]=(\alpha-1)e_{4},$ $[e_{1},e_{3},e_{2}]=\alpha e_{4}.$
$\mathfrak{B}_{06}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{3}]=e_{4},$
$[e_{2},e_{3},e_{1}]=(\alpha-1)e_{4},$ $[e_{1},e_{3},e_{2}]=\alpha e_{4}.$
$\mathfrak{B}_{07}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=\alpha e_{4},$ $[e_{1},e_{2},e_{1}]=e_{4},$ $[e_{1},e_{2},e_{3}]=e_{4},$
$[e_{2},e_{3},e_{1}]=e_{4},$ $[e_{1},e_{3},e_{2}]=2e_{4}.$
$\mathfrak{B}_{08}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=\alpha e_{4},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{4},$
$[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=e_{4},$ $[e_{1},e_{3},e_{2}]=2e_{4}.$
$\mathfrak{B}_{09}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{1},e_{3},e_{2}]=e_{4}.$
$\mathfrak{B}_{10}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$
$[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{1},e_{3},e_{2}]=e_{4}.$
Definition 9.
[…]
@@
$\mathfrak{L}_{1}$ : $\{e_{1},e_{2}\}=e_{3}.$
$\mathfrak{L}_{2}^{\alpha}$ : $[e_{1},e_{2}]=e_{3},$$\{e_{1},e_{2}\}=\alpha e_{3}.$
Theorem 11 (see [29]).
@@
${L}_{01}$ : $\{e_{1},e_{2}\}=e_{3}.$
${L}_{02}^{\alpha}$ : $[e_{1},e_{2}]=e_{3},$$\{e_{1},e_{2}\}=\alpha e_{3}.$
${L}_{03}$ :::: $[e_{2},e_{3}]=e_{4},$$\{e_{1},e_{3}\}=e_{4}.$
${L}_{04}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3}]=e_{4},$$\{e_{1},e_{2}\}=e_{4}.$
${L}_{05}^{\alpha}$ : $[e_{2},e_{3}]=\alpha e_{4},$$\{e_{1},e_{2}\}=e_{3},$$\{e_{2},e_{3}\}=e_{4}.$
${L}_{06}$ : $[e_{1},e_{3}]=e_{4},$$\{e_{1},e_{2}\}=e_{3},$$\{e_{2},e_{3}\}=e_{4}.$
${L}_{07}$ :::: $[e_{1},e_{2}]=e_{4},$$\{e_{1},e_{2}\}=e_{3},$$\{e_{2},e_{3}\}=e_{4}.$
${L}_{08}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$$\{e_{1},e_{2}\}=\alpha e_{3},$$\{e_{2},e_{3}\}=e_{4}.$
${L}_{09}^{\alpha,\beta}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3}]=e_{4},$$\{e_{1},e_{2}\}=\alpha e_{3},$$\{e_{2},e_{3}\}=\beta e_{4}.$
${L}_{10}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3}]=e_{4},$$\{e_{1},e_{2}\}=\alpha e_{3},$$\{e_{1},e_{3}\}=e_{4}.$
@@ 2.1 Preliminaries: geometric classification
Given a complex vector space ${\mathbb{V}}$ of dimension $n$, the set of bilinear and triliners maps
$\textrm{Bil}({\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})$cong $\textrm{Hom}({\mathbb{V}}^{\otimes 2},{\mathbb{V}})$cong $({\mathbb{V}}^{*})^{\otimes 2}\otimes{\mathbb{V}},$
$\textrm{Tril}({\mathbb{V}}\times{\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})$cong $\textrm{Hom}({\mathbb{V}}^{\otimes 3},{\mathbb{V}})$cong $({\mathbb{V}}^{*})^{\otimes 3}\otimes{\mathbb{V}}$
are vector spaces of dimension $n^{3}$ and $n^{4}$, respectively. The set of pairs of bilinear and triliners maps
@@
which is a vector space of dimension $n^{4}+n^{3}$. This vector space has the structure of the affine space $\mathbb{C}^{n^{4}+n^{3}}$ in the following sense: fixed a basis $e_{1},\ldots,e_{n}$ of ${\mathbb{V}}$, then any pair $(\mu,\mu^{\prime})\in\textrm{Bil}({\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})\oplus\textrm{Tril}({\mathbb{V}}\times{\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})$ is determined by some parameters $c_{i,j}^{k},c_{i,j,k}^{l}\in\mathbb{C}$, called structural constants, such that
$\mu(e_{i},e_{j})=\sum_{k=1}^{n}c_{i,j}^{k}e_{k}\textrm{ and }\mu^{\prime}(e_{i},e_{j},e_{k})=\sum_{l=1}^{n}c_{i,j,k}^{l}e_{t}$
which corresponds to a point in the affine space $\mathbb{C}^{n^{4}+n^{3}}$. Then a subset $\mathcal{S}$ of $\textrm{Bil}({\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})\oplus\textrm{Tril}({\mathbb{V}}\times{\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})$ corresponds to an algebraic variety, i.e., a Zariski closed set, if there are some polynomial equations in variables $c_{i,j}^{k},c_{i,j,k}^{l}$ with zero locus equal to the set of structural constants of the pairs in $\mathcal{S}$.
Given the identities defining Lie-Yamaguti (Bol) algebras, we can obtain a set of polynomial equations in the variables $c_{i,j}^{k},c_{i,j,k}^{l}$. This class of $n$-dimensional Lie-Yamaguti (Bol) algebras is a variety. Denote it by $\mathcal{T}_{n}$. Now, consider the following action of $\rm{GL}({\mathbb{V}})$ on ${\mathcal{T}}_{n}$:
@@
for $g\in\rm{GL}({\mathbb{V}})$, $(\mu,\mu^{\prime})\in\mathcal{T}_{n}$ and for any $x,y,z\in{\mathbb{V}}$. Observe that the $\textrm{GL}({\mathbb{V}})$-orbit of $(\mu,\mu^{\prime})$, denoted $O((\mu,\mu^{\prime}))$, contains all the structural constants of the pairs isomorphic to the Lie-Yamaguti (Bol) algebra with structural constants $(\mu,\mu^{\prime})$.
@@
Definition 12.
Let ${\rm T}$ and ${\rm T}^{\prime}$ be two $n$-dimensional Lie-Yamaguti (Bol) algebras corresponding to the variety $\mathcal{T}_{n}$ and $(\mu,\mu^{\prime}),(\lambda,\lambda^{\prime})\in\mathcal{T}_{n}$ be their representatives in the affine space, respectively. The algebra ${\rm T}$ is said to degenerate to ${\rm T}^{\prime}$, and we write ${\rm T}\to{\rm T}^{\prime}$, if $(\lambda,\lambda^{\prime})\in\overline{O((\mu,\mu^{\prime}))}$. If ${\rm T}\not\cong{\rm T}^{\prime}$, then we call it a proper degeneration. Conversely, if $(\lambda,\lambda^{\prime})\not\in\overline{O((\mu,\mu^{\prime}))}$ then we say that ${{\rm T}}$ does not degenerate to ${{\rm T}}^{\prime}$ and we write ${{\rm T}}\not\to{{\rm T}}^{\prime}$.
@@
Let ${{\rm T}}(*)=\{{{\rm T}}(\alpha):{\alpha\in I}\}$ be a family of $n$-dimensional Lie-Yamaguti (Bol) algebras corresponding to ${\mathcal{T}}_{n}$ and let ${{\rm T}}^{\prime}$ be another $n$-dimensional Lie-Yamaguti (Bol) algebra. Suppose that ${{\rm T}}(\alpha)$ is represented by the structure $(\mu(\alpha),\mu^{\prime}(\alpha))\in{\mathcal{T}}_{n}$ for $\alpha\in I$ and ${{\rm T}}^{\prime}$ is represented by the structure $(\lambda,\lambda^{\prime})\in{\mathcal{T}}_{n}$. We say that the family ${{\rm T}}(*)$ degenerates to ${{\rm T}}^{\prime}$, and write ${{\rm T}}(*)\to{{\rm T}}^{\prime}$, if $(\lambda,\lambda^{\prime})\in\overline{\{O((\mu(\alpha),\mu^{\prime}(\alpha)))\}_{\alpha\in I}}$. Conversely, if $(\lambda,\lambda^{\prime})\not\in\overline{\{O((\mu(\alpha),\mu^{\prime}(\alpha)))\}_{\alpha\in I}}$ then we call it a non-degeneration, and we write ${{\rm T}}(*)\not\to{{\rm T}}^{\prime}$.
@@
Observe that ${\rm T}^{\prime}$ corresponds to an irreducible component of $\mathcal{T}_{n}$ (more precisely, $\overline{{\rm T}^{\prime}}$ is an irreducible component) if and only if ${{\rm T}}\not\to{{\rm T}}^{\prime}$ for any $n$-dimensional Lie-Yamaguti (Bol) algebra ${\rm T}$ and ${{\rm T}(*)}\not\to{{\rm T}}^{\prime}$ for any parametric family of $n$-dimensional Lie-Yamaguti (Bol) algebras ${\rm T}(*)$. In this case, we will use the next ideas to prove that a particular algebra corresponds to an irreducible compotent. Firstly, since $\mathrm{dim}\,O((\mu,\mu^{\prime}))=n^{2}-\mathrm{dim}\,\mathfrak{Der}({\rm T})$, then if ${\rm T}\to{\rm T}^{\prime}$ and ${\rm T}\not\cong{\rm T}^{\prime}$, we have that $\mathrm{dim}\,\mathfrak{Der}({\rm T})<\mathrm{dim}\,\mathfrak{Der}({\rm T}^{\prime})$, where $\mathfrak{Der}({\rm T})$ denotes the Lie algebra of derivations of ${\rm T}$. Secondly, to prove degenerations, let ${{\rm T}}$ and ${{\rm T}}^{\prime}$ be two Lie-Yamaguti (Bol) algebras represented by the structures $(\mu,\mu^{\prime})$ and $(\lambda,\lambda^{\prime})$ from ${{\mathcal{T}}}_{n}$, respectively. Let $c_{i,j}^{k},c_{i,j,k}^{l}$ be the structure constants of $(\lambda,\lambda^{\prime})$ in a basis $e_{1},\dots,e_{n}$ of ${\mathbb{V}}$. If there exist $n^{2}$ maps $a_{i}^{j}(t):\mathbb{C}^{*}\to\mathbb{C}$ such that $E_{i}(t)=\sum_{j=1}^{n}a_{i}^{j}(t)e_{j}$ ($1\leq i\leq n$) form a basis of ${\mathbb{V}}$ for any $t\in\mathbb{C}^{*}$ and the structure constants $c_{i,j}^{k}(t),c_{i,j,k}^{l}(t)$ of $(\mu,\mu^{\prime})$ in the basis $E_{1}(t),\dots,E_{n}(t)$ satisfy $\lim\limits_{t\to 0}c_{i,j}^{k}(t)=c_{i,j}^{k}$ and $\lim\limits_{t\to 0}c_{i,j,k}^{l}(t)=c_{i,j,k}^{l}$, then ${{\rm T}}\to{{\rm T}}^{\prime}$. In this case, $E_{1}(t),\dots,E_{n}(t)$ is called a parametrized basis for ${{\rm T}}\to{{\rm T}}^{\prime}$. If $E_{e_{1}}^{t},E_{e_{2}}^{t},E_{e_{3}}^{t},E_{e_{4}}^{t}$ is a parametric basis for ${\bf A}\to{\bf B}$, then we denote a degeneration by ${\bf A}\xrightarrow{(E_{e_{1}}^{t},E_{e_{2}}^{t},E_{e_{3}}^{t},E_{e_{4}}^{t})}{\bf B}$.
Thirdly, to prove non-degenerations we may use a remark that follows from this lemma, see [4] and the references therein.
Lemma 14.
Consider two Lie-Yamaguti (Bol) algebras ${\rm T}$ and ${\rm T}^{\prime}$. Suppose ${\rm T}\to{\rm T}^{\prime}$. Let ${\mathcal{Z}}$ be a Zariski closed set in ${\mathcal{T}}_{n}$ that is stable by the action of the invertible upper (lower) triangular matrices. Then if there is a representation $(\mu,\mu^{\prime})$ of ${\rm T}$ in ${\mathcal{Z}}$, then there is a representation $(\lambda,\lambda^{\prime})$ of ${\rm T}^{\prime}$ in ${\mathcal{Z}}$.
To apply this lemma, we will give the explicit definition of the appropriate stable Zariski closed ${\mathcal{Z}}$ in terms of the variables $c_{i,j}^{k},c_{i,j,k}^{l}$ in each case. For clarity, we assume by convention that $c_{i,j}^{k}=0$ (resp. $c_{i,j,k}^{l}=0$) if $c_{i,j}^{k}$ (resp. $c_{i,j,k}^{l}$) is not explicitly mentioned on the definition of ${\mathcal{Z}}$.
Remark 15.
Moreover, let ${{\rm T}}$ and ${{\rm T}}^{\prime}$ be two Lie-Yamaguti (Bol) algebras represented by the structures $(\mu,\mu^{\prime})$ and $(\lambda,\lambda^{\prime})$ from ${{\rm T}}_{n}$. Suppose ${\rm T}\to{\rm T}^{\prime}$. Then if $\mu,\lambda$ represents algebras ${\rm T}_{0},{\rm T}^{\prime}_{0}$ in the affine space $\mathbb{C}^{n^{3}}$ and $\mu^{\prime},\lambda^{\prime}$ represents algebras ${\rm T}_{1},{\rm T}^{\prime}_{1}$ in the affine space $\mathbb{C}^{n^{4}}$ of algebras with a single multiplication, respectively, we have ${\rm T}_{0}\to{\rm T}^{\prime}_{0}$ and ${\rm T}_{1}\to{\rm T}^{\prime}_{1}$. So for example, $(0,\mu)$ can not degenerate in $(\lambda,0)$ unless $\lambda=0$.

### v2

@@ 1 Preliminaries: the algebraic classification
Identities $({\rm B2})$$({\rm B4})$ mean that a Bol algebra is a Lie triple system with respect to the trilinear operation [,,][-,-,-]. Thus, Lie–Yamaguti algebras and Bol algebras are generalizations of the Lie triple systems.
Theorem 1.8 (see [2]).
Let $\mathfrak{B}$ be a 4-dimensional complex nilpotent Bol algebra. Then $\mathfrak{B}$ is isomorphic to one of the following Lie–Yamaguti algebras $\mathcal{L}_{01},\dots,\mathcal{L}_{41},$ $\mathcal{L}_{45}^{\alpha},\dots,\mathcal{L}_{52}$ or one of the following mutually non-isomorphic algebras:
$\mathfrak{B}_{01}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{2},e_{3}]=e_{4}, [$$e_{2},e_{3},e_{1}]=-\frac{1}{2}e_{4},$
$e_{1},e_{3},e_{2}]=\frac{1}{2}e_{4}, [$$e_{2},e_{3},e_{2}]=e_{4}.$
$\mathfrak{B}_{02}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{3}]=e_{4}, [$$e_{2},e_{3},e_{1}]=-\frac{1}{2}e_{4},$
$e_{1},e_{3},e_{2}]=\frac{1}{2}e_{4}, [$$e_{2},e_{3},e_{2}]=e_{4}.$
$\mathfrak{B}_{03}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{2},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{3}]=e_{4}, [$$e_{2},e_{3},e_{1}]=-\frac{1}{2}e_{4},$
$e_{1},e_{3},e_{2}]=\frac{1}{2}e_{4}, [$$e_{2},e_{3},e_{2}]=e_{4}.$
$\mathfrak{B}_{04}^{\alpha}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{2},e_{3}]=e_{4}, [$$e_{2},e_{3},e_{1}]=(\alpha-1)e_{4}, [$$e_{1},e_{3},e_{2}]=\alpha e_{4}.$
$\mathfrak{B}_{05}^{\alpha}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{3}]=e_{4},$
$e_{2},e_{3},e_{1}]=(\alpha-1)e_{4}, [$$e_{1},e_{3},e_{2}]=\alpha e_{4}.$
$\mathfrak{B}_{06}^{\alpha}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{3}]=e_{4}, [$$e_{2},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{3}]=e_{4},$
$e_{2},e_{3},e_{1}]=(\alpha-1)e_{4}, [$$e_{1},e_{3},e_{2}]=\alpha e_{4}.$
$\mathfrak{B}_{07}^{\alpha}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{3}]=\alpha e_{4}, [$$e_{1},e_{2},e_{1}]=e_{4}, [$$e_{1},e_{2},e_{3}]=e_{4},$
$e_{2},e_{3},e_{1}]=e_{4}, [$$e_{1},e_{3},e_{2}]=2e_{4}.$
$\mathfrak{B}_{08}^{\alpha}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{3}]=\alpha e_{4}, [$$e_{2},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{4},$
$e_{1},e_{2},e_{3}]=e_{4}, [$$e_{2},e_{3},e_{1}]=e_{4}, [$$e_{1},e_{3},e_{2}]=2e_{4}.$
$\mathfrak{B}_{09}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{2},e_{3}]=e_{4}, [$$e_{1},e_{3},e_{2}]=e_{4}.$
$\mathfrak{B}_{10}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{3},$
$e_{1},e_{2},e_{3}]=e_{4}, [$$e_{1},e_{3},e_{2}]=e_{4}.$
Definition 1.9.
A compatible Lie algebra is vector space $\mathfrak{g}$ with two bilinear multiplications [,][-,-] and $\{-,-\}$, where $\mathfrak{g}_{1}=(\mathfrak{g},[-,-])$ and $\mathfrak{g}_{2}=(\mathfrak{g},\{-,-\})$ are Lie algebras and the two operations are required to satisfy the following identity
@@
$\mathfrak{L}_{1}$ :: $\{e_{1},e_{2}\}=e_{3}.$
$\mathfrak{L}_{2}^{\alpha}$ :: $e_{1},e_{2}]=e_{3},$ $\{e_{1},e_{2}\}=\alpha e_{3}.$
Theorem 1.11 (see [29]).
@@
${L}_{01}$ :: $\{e_{1},e_{2}\}=e_{3}.$
${L}_{02}^{\alpha}$ :: $e_{1},e_{2}]=e_{3},$ $\{e_{1},e_{2}\}=\alpha e_{3}.$
${L}_{03}$ :: $e_{2},e_{3}]=e_{4},$ $\{e_{1},e_{3}\}=e_{4}.$
${L}_{04}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{2},e_{3}]=e_{4},$ $\{e_{1},e_{2}\}=e_{4}.$
${L}_{05}^{\alpha}$ :: $e_{2},e_{3}]=\alpha e_{4},$ $\{e_{1},e_{2}\}=e_{3},$ $\{e_{2},e_{3}\}=e_{4}.$
${L}_{06}$ :: $e_{1},e_{3}]=e_{4},$ $\{e_{1},e_{2}\}=e_{3},$ $\{e_{2},e_{3}\}=e_{4}.$
${L}_{07}$ :: $e_{1},e_{2}]=e_{4},$ $\{e_{1},e_{2}\}=e_{3},$ $\{e_{2},e_{3}\}=e_{4}.$
${L}_{08}^{\alpha}$ :: $e_{1},e_{2}]=e_{3},$ $\{e_{1},e_{2}\}=\alpha e_{3},$ $\{e_{2},e_{3}\}=e_{4}.$
${L}_{09}^{\alpha,\beta}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{2},e_{3}]=e_{4},$ $\{e_{1},e_{2}\}=\alpha e_{3},$ $\{e_{2},e_{3}\}=\beta e_{4}.$
${L}_{10}^{\alpha}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{2},e_{3}]=e_{4},$ $\{e_{1},e_{2}\}=\alpha e_{3},$ $\{e_{1},e_{3}\}=e_{4}.$
@@ 2.1 Preliminaries: geometric classification
Given a complex vector space $\mathbb{V}$} of dimension $n$, the set of bilinear and trilinear maps
$\textrm{Bil}({\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})$ $\cong$ $\textrm{Hom}({\mathbb{V}}^{\otimes 2},{\mathbb{V}})$ $\cong$ $({\mathbb{V}}^{*})^{\otimes 2}\otimes{\mathbb{V}},$
$\textrm{Tril}({\mathbb{V}}\times{\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})$ $\cong$ $\textrm{Hom}({\mathbb{V}}^{\otimes 3},{\mathbb{V}})$ $\cong$ $({\mathbb{V}}^{*})^{\otimes 3}\otimes{\mathbb{V}}$
are vector spaces of dimension $n^{3}$ and $n^{4}$, respectively. The set of pairs of bilinear and trilinear maps
@@
is a vector space of dimension $n^{4}+n^{3}$. This vector space has the structure of the affine space $\mathbb{C}^{n^{4}+n^{3}}$ in the following sense: fixed a basis $e_{1},\ldots,e_{n}$ of $\mathbb{V}$}, any pair $(\mu,\mu^{\prime})\in\textrm{Bil}({\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})\oplus\textrm{Tril}({\mathbb{V}}\times{\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})$ is determined by some parameters $c_{i,j}^{k},c_{i,j,k}^{l}\in\mathbb{C}$, called structural constants, such that
$\mu(e_{i},e_{j})=\sum_{k=1}^{n}c_{i,j}^{k}e_{k}\textrm{ and }\mu^{\prime}(e_{i},e_{j},e_{k})=\sum_{l=1}^{n}c_{i,j,k}^{l}e_{l}$
which correspond to a point in the affine space $\mathbb{C}^{n^{4}+n^{3}}$. Then a subset $\mathcal{S}$ of $\textrm{Bil}({\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})\oplus\textrm{Tril}({\mathbb{V}}\times{\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})$ corresponds to an algebraic variety, i.e., a Zariski closed set, if there are some polynomial equations in variables $c_{i,j}^{k},c_{i,j,k}^{l}$ with zero locus equal to the set of structural constants of the pairs in $\mathcal{S}$.
Given the identities defining Lie–Yamaguti (Bol) algebras, we can obtain a set of polynomial equations in the variables $c_{i,j}^{k},c_{i,j,k}^{l}$. This class of $n$-dimensional Lie–Yamaguti (Bol) algebras is a variety. Denote it by $\mathcal{T}_{n}$. Now, consider the following action of $\rm{GL}({\mathbb{V}})$ on ${\mathcal{T}}_{n}$:
@@
for $g\in\rm{GL}({\mathbb{V}})$, $(\mu,\mu^{\prime})\in\mathcal{T}_{n}$ and for any $x,y,z\in{\mathbb{V}}$. Observe that the $\textrm{GL}({\mathbb{V}})$-orbit of $(\mu,\mu^{\prime})$, denoted $O((\mu,\mu^{\prime}))$, contains all the structural constants of the pairs isomorphic to the Lie–Yamaguti (Bol) algebra with structural constants $(\mu,\mu^{\prime})$.
@@
Definition 2.1.
Let $\rm T$} and $\rm T$$^{\prime}$ be two $n$-dimensional Lie–Yamaguti (Bol) algebras corresponding to the variety $\mathcal{T}_{n}$, and let $(\mu,\mu^{\prime}),(\lambda,\lambda^{\prime})\in\mathcal{T}_{n}$ be their representatives in the affine space, respectively. The algebra $\rm T$} is said to degenerate to $\rm T$$^{\prime}$, and we write ${\rm T}\to{\rm T}^{\prime}$, if $(\lambda,\lambda^{\prime})\in\overline{O((\mu,\mu^{\prime}))}$. If ${\rm T}\not\cong{\rm T}^{\prime}$, then we call it a proper degeneration. Conversely, if $(\lambda,\lambda^{\prime})\not\in\overline{O((\mu,\mu^{\prime}))}$ then we say that $\rm T$}} does not degenerate to $\rm T$$^{\prime}$ and we write ${{\rm T}}\not\to{{\rm T}}^{\prime}$.
@@
Let ${{\rm T}}(*)=\{{{\rm T}}(\alpha):{\alpha\in I}\}$ be a family of $n$-dimensional Lie–Yamaguti (Bol) algebras corresponding to ${\mathcal{T}}_{n}$ and let $\rm T$$^{\prime}$ be another $n$-dimensional Lie–Yamaguti (Bol) algebra. Suppose that ${{\rm T}}(\alpha)$ is represented by the structure $(\mu(\alpha),\mu^{\prime}(\alpha))\in{\mathcal{T}}_{n}$ for $\alpha\in I$ and $\rm T$$^{\prime}$ is represented by the structure $(\lambda,\lambda^{\prime})\in{\mathcal{T}}_{n}$. We say that the family ${{\rm T}}(*)$ degenerates to $\rm T$$^{\prime}$, and write ${{\rm T}}(*)\to{{\rm T}}^{\prime}$, if $(\lambda,\lambda^{\prime})\in\overline{\{O((\mu(\alpha),\mu^{\prime}(\alpha)))\}_{\alpha\in I}}$. Conversely, if $(\lambda,\lambda^{\prime})\not\in\overline{\{O((\mu(\alpha),\mu^{\prime}(\alpha)))\}_{\alpha\in I}}$ then we call it a non-degeneration, and we write ${{\rm T}}(*)\not\to{{\rm T}}^{\prime}$.
@@
Observe that $\rm T$$^{\prime}$ corresponds to an irreducible component of $\mathcal{T}_{n}$ (more precisely, $\overline{{\rm T}^{\prime}}$ is an irreducible component) if and only if ${{\rm T}}\not\to{{\rm T}}^{\prime}$ for any $n$-dimensional Lie–Yamaguti (Bol) algebra $\rm T$} and ${{\rm T}(*)}\not\to{{\rm T}}^{\prime}$ for any parametric family of $n$-dimensional Lie–Yamaguti (Bol) algebras ${\rm T}(*)$. In this case, we will use the next ideas to prove that a particular algebra corresponds to an irreducible component. Firstly, since $\mathrm{dim}\,O((\mu,\mu^{\prime}))=n^{2}-\mathrm{dim}\,\mathfrak{Der}({\rm T})$, it follows that if ${\rm T}\to{\rm T}^{\prime}$ and ${\rm T}\not\cong{\rm T}^{\prime}$, then $\mathrm{dim}\,\mathfrak{Der}({\rm T})<\mathrm{dim}\,\mathfrak{Der}({\rm T}^{\prime})$, where $\mathfrak{Der}({\rm T})$ denotes the Lie algebra of derivations of $\rm T$}.